ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 82B21, 82B05, 82B31 Shchepan’uk, Gennady, Institute for Mathematics of Ukrainian National Science Academy, Kyiv, Ukraine; Kondratiev, Yuri, Bonn University, Germany; Minlos, Robert, Institute for Information Transmission Problems of Russian National Science Academy, Moscow, Russia; Röckner, Michael, Bielefeld University, Germany Almost L_{2}-exponential mixing for classical continuous systems Using the method of cluster expansions, we prove the exponential L_{2+ε}- mixing (ε is arbitrary small and positive) for the Gibbs measures corresponding to dilute continuous systems of particles interacting via a stable, exponentially decreasing, pair potential.